determine the break even points of the profit function P(x)=-2x^2+7x+8, where x is the number of dirt bikes produced, in thousands
I tried to find the discriminant through b^2-4ac but i dont think that is what i am supposed to do.
I also tried to solve for the profit=0 so 0=-2x^2+7x+8, but i am unsure about the algebra

Answer by ewatrrr(24785)

 
Hi,
P(x)=-2x^2+7x+8
= -2(x - 7/4)^2 + 49/8 + 64/8
0 = -2(x - 7/4)^2 + 113/8
113/16 = (x-7/4)^2
7/4 ± sqrt(113)/4 = x
1.75 ± 2.6575 = x